Scot and Vidia, married taxpayers, earn $218,000 in taxable income and $5,000 in interest from an investment in City of Tampa bonds (Use the U.S. tax rate schedule for married filing jointly). Required a. If Scot and Vidia earn an additional $99,750 of taxable income, what is their marginal tax rate on this income? b. What is their marginal rate if, instead, they report an additional $99,750 in deductions? For all requirements, do not round intermediate calculations. Round your answer to 2 decimal places.) Marginal tax rate Marginal tax rate
Answer #1

Hi

Let me know in case you face any issue:

As per 2018, tax schedule: Married Filing Jointly: Solution: 24.22% Answer a] Explanation: Marginal tax rate Change in Tax/ Change in Taxable income Tax on old income (28179) + (218000-165000)*24% = Tax on New income (64179) + (317750-315000)*32% Change in Tax Change in Taxable income Marginal Tax Rate $ 40,899.00 65,059.00 $24,160.00 $99,750.00 24.22% Answer b) Explanation: 23.06% Marginal tax rate = Change in Tax / Change in Taxable income Tax on old income (28179) + (218000-165000)*24% = Tax on New income ( 7) + (118250-77400)*22% = Change in Tax Change in Taxable income Marginal Tax Rate $ 40,899.00 $ 17,894.00 $-23,005.00 $-99,750.00 23.06%